a lorry starts from rest and moves eith a constant acceleration of 8ms-2 . compute its velocity and the distance travelled after 4s has ellapsed
1
Expert's answer
2020-10-30T13:19:36-0400

The distance travelled by a body with constant acceleration (and zero initial speed) is:


"d = \dfrac{at^2}{2}"

where "a = 8m\/s^2" is its acceleration, "t = 4s" is ellapsed time. Thus, obtain:


"d = \dfrac{at^2}{2} = \dfrac{8\cdot 4^2}{2} = 64m"

The velocity is:


"v = at = 8m/s^2\cdot 4s = 32 m/s"

Answer. 32 m/s, 64m.